OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (3, 0) and a directrix at x = -3.

OpenStudy (jdoe0001):

so... any ideas on where the vertex may be?

OpenStudy (anonymous):

it would be at 3,0 right?

OpenStudy (jdoe0001):

thought that was the focus point?

OpenStudy (anonymous):

oh then not really no

OpenStudy (anonymous):

i know i need to use the equation 4py=x^2 but im not sure how

OpenStudy (jdoe0001):

well, notice, the directrix is x = -3, meaning is vertical one also bear in mind that, the vertex is right between the directrix, and the focus point thus 

OpenStudy (jdoe0001):

well. I should have said, the vertex is half-way between the directrix and focus point so, the focus point is a "p" distance from the vertex, and so is the directrix, a same "p" distance from it so, where would that put it anyway?

OpenStudy (anonymous):

on 0,0

OpenStudy (jdoe0001):

yeap notice the distance "p" from the focus point and the directrix, is 3, thus "p" is 3 or p is +3, because the parabola is opening towards the right so, now we know the vertex coordinates and "p" is +3 the parabola is opening horizontally, thus the squared variable is the "y" then thus \(\begin{array}{llll} (y-{\color{blue}{ k}})^2=4{\color{purple}{ p}}(x-{\color{brown}{ h}}) \\ (x-{\color{brown}{ h}})^2=4{\color{purple}{ p}}(y-{\color{blue}{ k}})\\ \end{array} \qquad \begin{array}{llll} vertex\ ({\color{brown}{ h}},{\color{blue}{ k}})\\ {\color{purple}{ p}}=\textit{distance from vertex to }\\ \qquad \textit{ focus or directrix} \end{array} \\ \quad \\ \quad \\ (y-{\color{blue}{ 0}})^2=4{\color{purple}{ (3)}}(x-{\color{brown}{ 0}})\)